OpenStudy (anonymous):

Given the sequence -1, 2, -4, 8, ..., identify the tenth term. Show your work.

OpenStudy (alexandervonhumboldt2):

look: 2/-1=-2 -4/2=-2 8/-4=-2 any ideas?

OpenStudy (alexandervonhumboldt2):

so the common ratio is -2

OpenStudy (anonymous):

But what formula will help me find the 10th ?

OpenStudy (alexandervonhumboldt2):

To find any term of an arithmetic sequence: \[a _{n}=a _{1}+\left( n-1 \right)*d\] where \[a _{1}\] is the first term of the sequence, d is the common difference, n is the number of the term to find.

OpenStudy (idku):

this is a geometric series, alexnader

OpenStudy (idku):

you are incorrect

OpenStudy (alexandervonhumboldt2):

woops thx @idku let me give another formula

OpenStudy (alexandervonhumboldt2):

To find any term of a geometric sequence: where \[a _{1}\] is the first term of the sequence, r is the common ratio, n is the number of the term to find. \[a _{n}=a _{1}*r^\left( n-1 \right)\]

OpenStudy (alexandervonhumboldt2):

thx @idku for correcting me.